Skip to Navigation (at bottom of page). (at bottom of page).


Meeting held on December 21, 2005.

The testing of spare equipment has reached a stage where the few remaining items (such as power supplies; PCI cards and motherboards) will only be tested as required.

Two monitors have been sold and the income from those used to purchase 4 plastic storage bins for the spare equipment. In addition one member has been given a monitor and soundcard.

A few power spikes have occurred recently but didn't create any problems.

A number of calls from members for information to resolve problems with their computers were received but only two machines actually required any hands on help in the past few weeks - both required major makeovers - thorough testing of components; replacement of faulty components; complete backups of all files on disks and reinstallation of OS and other software. A further 2 machines have been booked in for the New Year.


Meeting held on November 30, 2005.

The month started quietly but it wasn't long before that changed. This has been a very busy month for the Treasurer - repairing and/or upgrading member's machines and answering phone enquiries regarding installation and use of particular software. The network has also suffered a few power spikes although the UPS handled them all with the exception of an extremely strong one on Monday night which caused some of the modems to disconnect and at least one of the computers to reboot. However on Monday the most serious problem occurred which resulted in the network being unreachable for about 5 hours - from 12pm to 5pm....a routing stuff-up which caused looping between Telstra's and Ozemail's servers and took SiS 2 to 3 hours to resolve on our behalf once we ascertained that the problem wasn't at this end or due to work at the local exchange.

The region's spare equipment is currently being tested and discarded if too old or found faulty. The region currently has a number of surplus old keyboards and monitors which tested OK. These will be offered as spares to any member who'd like to take them off our hands to alleviate the growing shortage of storage space at the hub.

On Sunday Craig delivered the remaining ex-Wollongong gear to us. Unfortunately the computer can't be tested as yet as it is missing the 512MB of 184 pin DDR SDRAM that it contained when it was originally purchased and the region doesn't have any of that type at all.


Meeting held on October 26, 2005.

ARM Meeting held.
Heavy rain caused some problems with phone lines and ADSL link
Link was down for a few hours.
Having problems with the newsgroup feed.
Webpages updated.


Meeting held on September 21, 2005.

After a serious hacker attack it was decided to upgrade server hardware and software with added security enhancements.


Meeting held on August 17, 2005.

Mostly quiet for the last month, apart from problems with Bombadil which will be replaced soon.

Membership steady.

That's about it (short, isn't it?).


No MC Meeting held this month.


Meeting held on June 20, 2005.

Chris is still waiting on an adapter to complete the setup of the wireless access point.

The latest draft of the Web page re-design displays well in all but the earliest browser versions and will; be implemented soon.

Current balance $1,070. Forecast profit of at least $185 indicates that Hunter will be the only APANA region to record a profit for this financial year.

Membership currently totals 43 members - 4 more than last meeting.

7 modems and 2 hard-drives were received this week. Unfortunately 1 hard-drive was found to be 'dead'. The former Wollongong computer is expected to be delivered here within the next few months by a Sydney member on a planned trip to Newcastle.

Some dialup modems are failing to disconnect properly and have required resetting on a regular basis. Now that we have the ex-Wollongong modems onhand in future when a problem is discovered with a particular modem it will be replaced immediately.

Since at least July 1st, there has been a NNTP connection problem where the external network behaves as if there's ~50% loss of outgoing packets destined for port 119. No other port traffic is affected. Everything checks out on our side of the link (firewall, news server to and from, operating system, etc.) Forensic information is being discussed with SISgroup, however nothing has been heard back from them as yet.

Cray firewalled from accepting mail directly from outside, it now only taking/sending mail either from Giles or Bombadil, to reduce forged email.

The ADSL connection needing to be manually reset several times in the past month.

A GPL'd email virus scanner 'ClamAV' is currently in testing. It should be installable on Bombadil, Giles, and Cray.

WebMAIL can be trivially exploited to let anyone on the network read anyone's email!

Carolyn asked if the recent Debian update had been installed. Chris advised that all the network machines are set to automatically update when new releases become available.

Carolyn advised of the progress made in compiling a CD containing APANA and Hunter APANA information and software to be given to new members and stated that the current CD compilation only uses around 400MB of space, so there is plenty of room to include other software.


Meeting held on May 18, 2005.

Chris successfully set up a wireless access point, and is waiting on hardware to make improvements. He has also set up a station identifier.

No progress report on Peter's Sys Admin page, probably because web page re-design is still incomplete.

Web page re-design is still in progress. Colour problems need to be resolved before the re-design is finalised.

Hunter APANA have enough spare parts to make three whole PCs! A Pentium board has been purchased and is slated for Bombadil. Carolyn has accepted a donated PC.

There have been SPAM e-mail attacks , and AOL was refusing mail temporarily. Sybas got trojanized but was later fixed. Chris has toughened the firewall. The new server change-over was causing active file renumbering in newsgroups. Froggy moved over to a permanent connection. At the start of March NetCentral had major dialling problems. The change-over to SiS took several days -- the new link is performing OK, though day traffic sometimes becomes overloaded.

New Rates

Chris proposed placing a "W" sign on the antenna mast at 21 Reservoir Rd, to indicate a "closed W.E.P. network". This is to be made from local materials at no expense.

Carolyn suggested creating and placing a sign for APANA in from of 21 Reservoir Rd. She will chase up quotes and present these at the next meeting.


Meeting held on April 13, 2005.

We lost a couple of members and we had to delay our last meeting due to illness so are holding it this Saturday.

There may be a fee reduction in the wind.


Meeting held on March 23, 2005.


Total members 39: 18 dialins 1 perms 3 orgs 9 uucp/mail-news 2 ethernet 6 APANA members only

Current balance is $1,147.24


The Hunter APANA ADSL link has been transferred from NetCentral to the SIS Group. However, due to bad coordination by NetCentral (they didn't advise SIS of this until the Monday) Hunter APANA were without the use of the ADSL link for a full weekend. There were further problems with the re-routing due to a misconfiguration, but this has now been fixed.

UPGRADES Chris expects to upgrade Bombadil soon with a PII/300 w/ 128MB motherboard.


The re-design of the Hunter APANA web pages is going slowly as the designer (yours truly) is now back at University, and has to give studies a priority. Current refit is changing the page column widths to be Em based, so that text can be easily resized by the viewer.


Meeting held February 16,2005.

cf: Regional Committee Minutes, 13/2/05.


Meeting held on January 19, 2005.

The region has been fairly quiet apart from one or two virus hits. There have been no responses to community newspaper ads we have placed.